OpenStudy (anonymous):

The number of lines formed by 4 points, no three of which are collinear, is ____. A.5 B.4 C.6 D.3

OpenStudy (anonymous):

Use the formula \[\frac{n(n-1)}{2}\]Which is the total number of connections between n points. So:\[\frac{4(3)}{2}=6\]

OpenStudy (anonymous):

12/2=6

OpenStudy (anonymous):

Yep.